# Geospatial analysis of toponyms in geotagged social media posts

### v1

@@ III.3 Location-independent model
This model essentially posits that the underlying mechanism of the occurrence of a word can be summarized by probability $p_{w,c}$ specific to each grid cell $c$. Importantly, it is based on the simplifying assumption that the correlations between occurrences of different words can be neglected.
@@
Figure 4A shows the distribution of occurrence of “Fukuoka” against $n_{\mathrm{all}}$ in empirical data and the expectation from the location-independent binomial model. We observe that the empirical distribution is not consistent with the model. In particular, the model does not reproduce the two-branch scaling behavior seen in the empirical data and exhibits a single scaling instead. We confirm the same kind of discrepancy for all the domestic toponyms we studied (see Fig S4 in Supporting Information for the full results).
@@
Figure 4: Relationship between $n_{w}$ and $n_{\mathrm{all}}$ in empirical and model distributions. (A) Scatter plots for different keywords. (B, C) Kernel density profiles of empirical and model distributions. In each panel, the red contours show the density profile obtained from the location-independent model (B) or the core-periphery model (C), fitted to the empirical occurrence pattern of each word, represented by the blue contours. As in Fig. 3D, the region inside each contour, from dark to light colors, contains 90.0%, 99.0%, and 99.9% of the data points, respectively. Note that the scatter plot of empirical data and its density profile for “Fukuoka” are identical to those in Fig 3D.
@@
The implication of this observation becomes apparent through juxtapositions against patterns for common nouns and foreign nouns. The location-independent model shows a close agreement with the empirical data for “wallet” (Fig 4B), implying that the word indeed occurs at a constant rate in every grid cell. In general, we find that the empirical data and the location-independent model are in fairly good agreement for many common nouns that refer to objects (e.g., “telephone”) or abstract concepts (e.g., “society”). For a comparison between the model and the empirical distributions for these words, see Fig S4 in Supporting Information. On the other hand, there is a large deviation between the empirical and model distributions for the word “airport” (Fig 4C), suggesting that the occurrence at a constant rate is not a shared characteristic among all common nouns. This can be explained by the fact that common words such as “airport” are often used in combination with toponyms (e.g. “Narita Airport”), and are therefore more likely to be used in specific places. Finally, for foreign toponyms, such as “Hawaii” shown in Fig 4D, we observe the single scaling pattern in both the empirical and model distributions, although the empirical distributions generally show slightly greater variance than the model. While these toponyms are not semantically associated with a specific place in the observed area, they may occur with higher probability around international airports from which people travel to the places these toponyms refer.
@@
Beyond visual inspection, the discrepancy between the location-independent model $P$ and the data can be quantified using relative entropy, also known as the Kullback–Leibler (KL) divergence. Intuitively, relative entropy measures the dissimilarity from one probability distribution to another. Here, we wish to quantify the dissimilarity of the empirical data from the model, each of which is a probability distribution on the total number of posts in each cell and the number of posts with word $w$ in each cell. Specifically, we employ a modified version of relative entropy, denoted by $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$, that takes into account the cells with at least one occurrence of $w$. We elaborate on this particular definition and provide the rational for our choice in Supporting Information S1.
@@
As shown in Fig 5, relative entropy $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$ clearly differentiates domestic toponyms and place nouns from foreign toponyms and common nouns without place connotations. All domestic toponyms except “Kanto” and all place nouns (“park”, “university”, “hotel”, “airport”, and “shrine”) are characterized by relatively large values of relative entropy, namely $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})>4$, implying that the empirical distribution is highly dissimilar from the location-independent binomial model for these words. Common nouns such as “trip” and “vegetable” are on the borderline ($D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})\simeq 4$); relative entropy for other common nouns and foreign toponyms are significantly smaller.
@@
III.4 Core-periphery model
@@
So far we have shown that the location-independent binomial model cannot account for the large variance and two-branch scaling behavior seen in empirical distribution for domestic toponyms. Taking a step towards reality while maintaining simplicity, we consider a location-dependent variant of the binomial model. Namely, we assume that the value of $p_{w,c}$ in Eq (2) can vary for different grid cell $c$.
If the probability $p_{w,c}$ that a post contains toponym $w$ is to be estimated independently for each grid cell $c$, its unbiased, maximum likelihood estimator would be equal to the occurrence ratio $\phi_{w,c}=n_{w,c}/n_{\mathrm{all},c}$. However, since this model has as many parameters as the number of grid cells, it provides little insight into the spatial patterns of toponym occurrence. Here we propose a simpler model that assumes the presence of a core, where the toponym occurs most frequently, and periphery, geographically distant grid cells with lower $p_{w,c}$. We naturally expect that the core indicates the geographical extent to which the toponym refers.
@@
Plotting the occurrence ratio $\phi_{w,c}$ against the geodesic distance $d_{w,c}$ between the center $O_{w}$ and cell $c$, we observe an overall decreasing trend with distance from the center (Fig 6A). In fact, the binned averages of $\phi_{w,c}$ as a function of distance from the center imply a power-law decay of the average occurrence ratio at locations far from the center. Moreover, by grouping toponyms according to administrative level, as shown in Fig 6B, one can see that the onsets of the power law differ according to the granularity of each toponym. For toponyms referring to higher-level administrative units, which are typically larger in area, the power-law decay starts at a larger distance, while the average occurrence ratios are relatively stable at smaller distances. Conversely, toponyms denoting small administrative units, such as wards, exhibit power-law behavior starting at short distances with no noticeable plateau regime.
@@
Figure 6: Core-periphery patterns of toponym occurrence. (A) Occurrence ratio $\phi_{w}$ of “Fukuoka” against distance $d_{w}$ from center $O_{w}$ (small black dots), overlaid with the average for each logarithmic bin (red circles). The solid and dashed lines represent the maximum likelihood fits of the location-independent and core-periphery binomial models. (B) Average occurrence ratio as a function of $d_{w}$ for all the domestic toponyms studied in this work. (C) Maximum likelihood estimator of the core-periphery model parameters for each toponym. We represent estimated radius $\hat{r}_{w}$ by bars colored according to the category of the toponym (lower axis) and estimated exponent $\hat{a}_{w}$ by gray circles (upper axis). (D) The fitted core-periphery model compared to the administrative/metropolitan area. The innermost circle in dark green represents the core boundary (distance $r_{w}$ from the center $O_{w}$) and the two outer circles in lighter green denote the distance at which the occurrence probability $p_{w,c}$ is equal to one half and one third of the probability in the core $p_{0,w}$, respectively. The purple shaded area indicates the administrative area of each prefecture (top row) and the metropolitan employment area of each city (bottom row).
@@
Motivated by these observations, our core-periphery model is formalized as follows. The occurrence probability of the binomial model is assumed to be constant within a certain distance from the center and decay as a power law outside of that range:
$p_{w,c}=\begin{cases}p_{0,w}&\text{for }d_{w,c}\leq r_{w},\\ p_{0,w}(d_{w,c}/r_{w})^{-a_{w}}&\text{for }d_{w,c}>r_{w}.\end{cases}$ (3)
@@
This model has three free parameters: $p_{0,w}$ is the occurrence probability in cells within a radius of $r_{w}$ from the center, and $a_{w}$ denotes the exponent of the power-law decay outside the range.
@@
We estimate the parameters of the space-dependent binomial model by maximizing the likelihood for the geographic distribution of each toponym. The fitted model shows, in general, a better agreement to the data in terms of the relationship between $n_{\mathrm{all}}$ and $n_{w}$, as shown in Fig 4B. For the results for all domestic toponyms studied in this work, see Fig S5 in Supporting Information. In particular, this model reproduces the two-branch scaling behavior of the empirical data. This significant improvement from the location-independent model is also quantitatively demonstrated by the decrease in Akaike information criterion (AIC) for all the toponyms studied in this work (Fig 7).
@@
The advantage of the core-periphery model is that it conforms to an intuitive interpretation: the radius $r_{w}$ can be seen as the extent of the area to which the toponym refers, i.e. the core. Outside this core area, users refer to the place less often, but the decrease in probability is gradual as a function of distance and slow enough to be modeled as a power law (compared to, e.g., an exponential decay). To verify this interpretation, we compare the estimated core (the area within the estimated radius $\hat{r}_{w}$ from the center $O_{w}$) of each toponym with the extent of the administrative unit or metropolitan area to which it refers (Fig 6D). For all toponyms, the model identifies the center within the area to which each toponym refers, although the sizes of the cores vary and do not necessarily coincide with the administrative boundaries. For many toponyms, the core area detected by the model is smaller than the administrative area. This may indicate that users are more likely to make geotagged posts with these toponyms when they are in the central city of a prefecture or in the central area of a city. For “Fukuoka” and “Sendai”, the core has a geographic scale similar to the metropolitan area, suggesting that the use of these two toponyms is aligned with the extent of the corresponding metropolitan area. In the case of “Sapporo”, the core is larger than the metropolitan area. This could indicate that users tend to associate the toponym with a larger area; however, it is also possible that this is because the estimated parameter represents one of the many local maxima in the likelihood landscape.
@@ IV Discussion and conclusion
Finally, we presented the core-periphery model, which assumes a location-dependent occurrence probability of toponyms. Despite its simplicity with only three fitting parameters, this model can reasonably reproduce the empirically observed geographic distributions of toponym occurrence. This implies the following: First, each toponym has a core, i.e., a geographic area in which the toponym occurs with the highest probability, which can be regarded as the area that users collectively identify with the toponym. This interpretation is supported by the fact that the core is larger for the names of regions than for the names of cities and wards. Second, outside this core, the occurrence probability decreases slowly with distance following a power law. This may indicate that human attention, perception, and representation of geographic space respond nonlinearly to distance. It is also reminiscent of Tobler’s first law of geography: “everything is related to everything else, but near things are more related than distant things” [20]. In this context, our findings can be seen as another example of distance decay behavior in human activities, alongside other aspects such as commuting [21, 22, 23], tourism [24, 25, 26], and crime [27, 28, 29]. Yet, only few works have examined distance decay in linguistic patterns [30, 18], and we believe that our analysis is the first to confirm it on a large scale.
@@
We also remark the possibility that geotagged posts on Twitter may not be an unbiased, representative sample of the language use of the general population. This issue can be divided into two questions: whether the geotagged Twitter users can be considered a good proxy for the population at large [31, 11, 32, 33], and whether the language use in geotagged Twitter posts is consistent with language use in other contexts [34, 35]. The effect of the first problem on our results is presumably relatively small compared to other work using geotagged Twitter posts to study language use, for two reasons: (1) we focus only on the content of the posts without correlating them with user demographics, and (2) the age and socioeconomic status of users are unlikely to significantly affect their use of toponyms. It is however possible that urban toponyms are overrepresented due to population biases, which could affect the geographic distribution of toponyms. In this work, we focused on relatively large cities and wards within them, but whether our findings generalize to the names of smaller towns and villages needs to be carefully examined in future research. The second question concerns the generalizability of our findings to the use of toponyms in other contexts. To answer this question, one needs to analyze data from other social media platforms or in other forms of communication. Although this is an important point, it is beyond the scope of our study and should be addressed in future work.

### v2

@@ III.2 Location-independent model
This model essentially posits that the underlying mechanism of the occurrence of a word can be summarized by probability $p_{w,c}$ specific to each grid cell $c$. Importantly, it relies on the simplifying assumption that the occurrence of a word in one post is independent of its occurrence in other posts, and that the dependence between the occurrence of different words within the same post can also be neglected.
@@
Figure 4A shows the distribution of occurrence of Fukuoka against $n_{\mathrm{all}}$ in empirical data and the expectation from the location-independent binomial model. We observe that the empirical distribution is not consistent with the model. In particular, the model does not reproduce the two-branch scaling behavior seen in the empirical data and exhibits a single scaling instead. We confirm the same kind of discrepancy for all the domestic toponyms we studied (see Fig S3 in Supporting Information for the full results).
@@
Figure 4: Relationship between $n_{w}$ and $n_{\mathrm{all}}$ in empirical and model distributions. (A) Scatter plots for different keywords. (B, C) Kernel density profiles of empirical and model distributions. In each panel, the red contours show the density profile obtained from the location-independent model (B) or the core-periphery model (C), fitted to the empirical occurrence pattern of each word, represented by the blue contours. As in Fig. 3D, the region inside each contour, from dark to light colors, contains 90.0%, 99.0%, and 99.9% of the data points, respectively. Note that the scatter plot of empirical data and its density profile for Fukuoka are identical to those in Fig 3D.
@@
The implication of this observation becomes apparent through juxtapositions against patterns for common nouns and foreign nouns. The location-independent model shows a close agreement with the empirical data for wallet (Fig 4B), implying that the word indeed occurs at a constant rate in every grid cell. In general, we find that the empirical data and the location-independent model are in fairly good agreement for many common nouns that refer to objects (e.g., telephone) or abstract concepts (e.g., society). For a comparison between the model and the empirical distributions for these words, see Fig S3 in Supporting Information. On the other hand, there is a large deviation between the empirical and model distributions for the word airport (Fig 4C), suggesting that the occurrence at a constant rate is not a shared characteristic among all common nouns. This can be explained by the fact that common words such as airport are often used in combination with toponyms (e.g. Narita Airport), and are therefore more likely to be used in specific places. Finally, for foreign toponyms, such as Hawaii shown in Fig 4D, we observe the single scaling pattern in both the empirical and model distributions, although the empirical distributions generally show slightly greater variance than the model. While these toponyms are not semantically associated with a specific place in the observed area, they may occur with higher probability around international airports from which people travel to the places these toponyms refer.
@@
Beyond visual inspection, the discrepancy between the location-independent model $P$ and the data can be quantified using relative entropy, also known as the Kullback–Leibler (KL) divergence. Intuitively, relative entropy measures the dissimilarity from one probability distribution to another. Here, we wish to quantify the dissimilarity of the empirical data from the model, each of which is a probability distribution on the total number of posts in each cell and the number of posts with word $w$ in each cell. Specifically, we employ a modified version of relative entropy, denoted by $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$, that takes into account the cells with at least one occurrence of $w$. We elaborate on this particular definition and provide the rationale for our choice in Supporting Information.
@@
As shown in Fig 5, relative entropy $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$ clearly differentiates domestic toponyms and place nouns from foreign toponyms and common nouns without place connotations. All domestic toponyms except Kanto and all place nouns (park, university, hotel, airport, and shrine) are characterized by relatively large values of relative entropy, namely $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})>4$, implying that the empirical distribution is highly dissimilar from the location-independent binomial model for these words. Common nouns such as trip and vegetable are on the borderline ($D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})\simeq 4$); relative entropy for other common nouns and foreign toponyms are significantly smaller.
@@
III.3 Core-periphery model
@@
So far, we have shown that the location-independent binomial model cannot reproduce the large variance and two-branch scaling behavior seen in the empirical distribution of domestic toponyms. Let us now take a step toward realism and discuss a more flexible modeling framework to account for the empirically observed geographic distribution of toponyms. We consider the location-dependent binomial model, that is, we assume that the occurrence probability $p_{w,c}$ in Eq (2) can vary for different grid cells.
A straightforward implementation of the location-dependent model would be to allow the occurrence probabilities for different cells to be independent of each other. In this case, the value of $p_{w,c}$ can be estimated individually for each $c$ and its unbiased maximum likelihood estimator would be equal to the occurrence ratio $\phi_{w,c}=n_{w,c}/n_{\mathrm{all},c}$. However, since this model has as many parameters as the number of grid cells, it provides little insight into the spatial patterns of toponym occurrence. What would be more interesting to us is a model that captures the empirical observation with a smaller number of parameters.
@@
Let us recall that the results in Fig 3 suggest that a post in grid cell $c$ is more likely to contain a toponym $w$ if $c$ is within the area that $w$ refers to. Indeed, the occurrence ratio $\phi_{w,c}$ plotted against the geodesic distance $d_{w,c}$ between the center $O_{w}$ and cell $c$ shows an overall decreasing trend (Fig 6A). The binned averages of $\phi_{w,c}$ imply that the occurrence probability decays as a power-law function of distance from the center, especially at long distances. Moreover, by grouping toponyms according to their administrative level as shown in Fig 6B, one can see that the onsets of the power law differ according to the granularity of each toponym. For toponyms referring to higher-level units, such as regions and prefectures, which are typically larger in area, the power-law decay starts at larger distances while the average occurrence ratios are relatively stable at smaller distances. Conversely, toponyms denoting small administrative units, such as wards, exhibit power-law behavior starting at short distances with no noticeable plateau regime.
@@
Figure 6: Core-periphery patterns of toponym occurrence. (A) Occurrence ratio $\phi_{w}$ of Fukuoka against distance $d_{w}$ from center $O_{w}$ (small black dots), overlaid with the average for each logarithmic bin (red circles). The solid and dashed lines represent the maximum likelihood fits of the location-independent and core-periphery binomial models. (B) Average occurrence ratio as a function of $d_{w}$ for all the domestic toponyms studied in this work. (C) Maximum likelihood estimator of the core-periphery model parameters for each toponym. We represent estimated radius $\hat{r}_{w}$ by bars colored according to the category of the toponym (lower axis) and estimated exponent $\hat{a}_{w}$ by gray circles (upper axis). The standard errors are omitted as they are too small to be meaningfully visualized. (D) The fitted core-periphery model compared to the administrative/metropolitan area. The innermost circles in dark green represent the core boundary (distance $r_{w}$ from the center $O_{w}$) and the two outer circles in lighter green denote the distance at which the occurrence probability $p_{w,c}$ is equal to one half and one third of the probability in the core $q_{w}$, respectively. The areas shaded in purple indicate the administrative area of each prefecture (top row) and the metropolitan employment area of each city (bottom row). Maps made with Natural Earth (https://www.naturalearthdata.com/).
@@
Motivated by these observations, we propose a simple location-dependent variant of the binomial model (Eq (2)) that assumes the presence of a core, the area in which the occurrence probability $p_{w,c}$ is high, and a periphery, grid cells that are geographically distant from the center and characterized by lower $p_{w,c}$. Namely, the occurrence probability is assumed to be constant within a certain distance from the center and to decrease as a power law outside of this range:
$p_{w,c}=\begin{cases}q_{w}&\text{for }d_{w,c}\leq r_{w},\\ q_{w}\left[\dfrac{d_{w,c}}{r_{w}}\right]^{-a_{w}}&\text{for }d_{w,c}>r_{w}.\end{cases}$ (3)
@@
This model has three free parameters: $q_{w}$ is the occurrence probability in cells within a radius of $r_{w}$ from the center, and $a_{w}$ denotes the exponent of the power-law decay outside the range. In Fig 6A, we visualize the behavior of this model as a function of distance from the center in comparison to that of the location-independent model.
For each toponym $w$, the parameter values of the core-periphery model are estimated by maximizing the log-likelihood function
$\mathcal{L}(q_{w},r_{w},a_{w})=\sum_{c}\log P(n_{w,c}\mid n_{\mathrm{all},c},\,p_{w,c}).$
[…]
@@
The fitted model shows, in general, a better agreement to the data in terms of the relationship between $n_{\mathrm{all}}$ and $n_{w}$, as shown in Fig 4B. For the results for all domestic toponyms studied in this work, see Fig S4 in Supporting Information. In particular, this model reproduces the two-branch scaling behavior of the empirical data. This significant improvement from the location-independent model is also evidenced quantitatively by the decrease in the Akaike Information Criterion (AIC) for all the toponyms studied in this work (Fig 7).
@@
The advantage of the core-periphery model is that it conforms to an intuitive interpretation: the radius $r_{w}$ can be seen as the extent of the area to which the toponym refers, i.e. the core. Outside this core area, users refer to the place less often, but the decrease in probability is gradual as a function of distance and slow enough to be modeled as a power law (compared to, e.g., an exponential decay). To verify this interpretation, we compare the estimated core (the area within the estimated radius $\hat{r}_{w}$ from the center $O_{w}$) of each toponym with the extent of the administrative unit or metropolitan area to which it refers (Fig 6D). For all toponyms, the model identifies the center within the area to which each toponym refers, although the sizes of the cores vary and do not necessarily coincide with the administrative boundaries. For many toponyms, the core area detected by the model is smaller than the administrative area. This may indicate that users are more likely to make geotagged posts with these toponyms when they are in the central city of a prefecture or in the central area of a city. For Fukuoka and Sendai, the core has a geographic scale similar to the metropolitan area, suggesting that the use of these two toponyms is aligned with the extent of the corresponding metropolitan area. In the case of Sapporo, the core is larger than the metropolitan area. This could indicate that users tend to associate the toponym with a larger area; however, it is also possible that this is because the estimated parameter represents one of the many local maxima in the likelihood landscape.
@@ IV Discussion and Conclusions
Finally, we presented the core-periphery model, which assumes a location-dependent occurrence probability of toponyms. Despite its simplicity with only three fitting parameters, this model can reasonably reproduce the empirically observed geographic distributions of toponym occurrence. This implies the following: First, each toponym has a core, i.e., a geographic area in which the toponym occurs with the highest probability, which can be regarded as the area that users collectively identify with the toponym. This interpretation is supported by the fact that the core is larger for the names of regions than for the names of cities and wards. Second, outside this core, the occurrence probability decreases slowly with distance following a power law.
Our findings may indicate that human attention, cognition, and representation of geographic space respond nonlinearly to distance [27]. It is reminiscent of Tobler’s first law of geography: “everything is related to everything else, but near things are more related than distant things” [28]. In this context, our findings can be seen as another example of the distance decay phenomenon, which has been observed in various aspects of human behavior such as commuting [29, 30, 31], tourism [32, 33, 34], and crime [35, 36, 37]. The concept of distance decay, and its more sophisticated formulation, the gravity model, has also been used in archaeology and history [38, 39], attesting to its universal applicability in describing human activities. Inspired by developments in geography, distance decay and gravity models have been adopted in linguistics to explain variations in pronunciation between different dialects [40, 41] and language evolution via lexical replacement [42]. However, unlike other language elements that are generally geography-neutral, each toponym is intrinsically associated with a specific geographic point or area. As such, our results present a unique variant of Tobler’s first law in language, one that cannot be characterized by autocorrelation or other similarity measures [43, 44].
@@
We also remark on the possibility that geotagged posts on Twitter may not be an unbiased, representative sample of the language use of the general population. This issue can be divided into two questions: whether the geotagged Twitter users can be considered a good proxy for the population at large [45, 13, 46, 47], and whether the language use in geotagged Twitter posts is consistent with language use in other contexts [48, 49]. The effect of the first problem on our results is presumably relatively small compared to other work using geotagged Twitter posts to study language use, for two reasons: (i) we focus only on the content of the posts without correlating them with user demographics, and (ii) the age and socioeconomic status of users are unlikely to significantly affect their use of toponyms. It is however possible that urban toponyms are overrepresented due to population biases, which could affect the geographic distribution of toponyms. In this work, we focused on relatively large cities and wards within them, but whether our findings generalize to the names of smaller towns and villages needs to be carefully examined in future research. The second question concerns the generalizability of our findings to the use of toponyms in other contexts. Indeed, prior research has shown that, for certain linguistic features, such as emoji and hashtags, the same user may exhibit different styles on different social platforms [50]. However, we expect that toponyms are unlikely to be strongly influenced by such contextual differences. Toponyms are a linguistic element that is relatively stable and resistant to change [51], and it is reasonable to assume that this stability stems from the low synchronic variability of toponym use in the population.
Lastly, we note that our models are phenomenological and do not account for the microscopic underpinnings behind the observation, such as the behavior of individual users or posts. Moreover, gaining a comprehensive understanding of the use of toponyms—how they reflect the interaction between people and environment, how they shape and reinforce people’s identity, and how they are affected by urban planning and place branding—would require historical, ecological, cultural, and economic analyses [52, 53, 54, 55, 56, 57, 58, 59]. These aspects are abstracted away in the present study, where our focus is to establish general empirical laws that govern the spatial distribution of toponyms. The simplified models we propose are designed to serve the purpose of quantitative, so-called extensive analysis [60]. Future work should aim to integrate these qualitative factors with our quantitative framework to investigate the dynamics underlying toponym usage and distribution.